# PartnerMAS: An LLM Hierarchical Multi-Agent Framework for Business Partner Selection on High-Dimensional Features

### v1

@@ 3.2 MAS Architecture Design
To tackle the challenge of partner selection, we introduce a hierarchical MAS framework called PartnerMAS (Figure 1 (C)). This design enables the system to decompose the high-dimensional task into manageable sub-tasks, each assessing tabular data from different dimensions: (i) Planner Agent (PAPA) that designs the evaluation strategy. (ii) Specialized Agents (SASA), $\{SA_{1},\dots,SA_{N}\}$, which are dynamically configured by the PAPA to execute this strategy based on their role definition and expertise. (iii) Supervisor Agent (SPASPA) that aggregates analyses to make a final decision.
@@
Planner Agent (PAPA). The primary role of PAPA is to interpret the high-level task context $\mathcal{Q}$ (Lead VC and target company profiles) and formulate an evaluation plan. It does so by analyzing $\mathcal{Q}$ with feature names to identify the most critical evaluation dimensions given the case description. The PAPA’s output is a set of NN agent configurations, $\{A_{1},\dots,A_{N}\}$, where each configuration $A_{i}$ contains a specific “profile” that guides the corresponding $SA_{i}$. It also generates “strategic guidance” to support the decision for the SPASPA. Formally, the output can be expressed as: $\{A_{1},\dots,A_{k}\},PA(\mathcal{Q},C_{sample})]$, a set of explicit instructions for the team of SASAs.
@@
Specialized Agent (SASA). Each SASA acts as a domain expert, tasked with evaluating the entire candidate pool $C$ from its specialized perspective. To manage the high dimensionality of the candidate feature vectors ($x_{j}\in\mathbb{R}^{n}$), the SASA first performs a feature selection step. Guided by its assigned $profile_{i}$, it identifies and focuses on a relevant feature subset. The agent’s core evaluation function $f_{i}$, driven by a backbone LLM, then directly produces a ranked shortlist of candidates. This output, $S^{\prime}_{i}$, is a list containing the top $k^{\prime}=\lceil m/3\rceil$ firms, where each entry includes the firm’s ID, its rank, and an alignment score, $score_{ij}\in[1,10]$.
@@
Supervisor Agent (SPASPA). The SPASPA is responsible for synthesizing the shortlists $\{F^{\prime}_{1},\dots,F^{\prime}_{N}\}$ from Specialized Agents into the final ranked list $F$. It mimics a human-led committee’s decision-making process by first establishing consensus and then resolving disagreements based on the strategic priorities of the deal. This is achieved in a two-step process: (i) Consensus Selection: The SPASPA first identifies candidates with broad support across SASA, by counting how many agents include them in their shortlists. This step can determine robust candidates that perform well across different dimensions: $F_{1}(c_{j})=\sum_{i=1}^{N}\mathbb{I}(c_{j}\in F^{\prime}_{i})$. (ii) Conflict Resolution: To fill the remaining slots in the shortlist, the SPASPA resolves disagreements among the SASAs. It first determines an importance weight $w_{i}$ for each agent based on the agent’s relevance to the specific deal. It then resolves the conflict for the remaining candidates by giving more weight to the opinions of more important SASAs. This allows the SPASPA to select candidates who excel in critical areas, even if they lack broad consensus: $F_{2}(c_{j})=\sum_{i=1}^{N}w_{i}\cdot\frac{1}{R_{i}(c_{j})}$. The final shortlist $F=[F_{1}(c_{j}),F_{2}(c_{j})]$.

### v2

@@ 3.2 MAS Architecture Design
To tackle the challenge of partner selection, we introduce a hierarchical MAS framework called PartnerMAS (Figure 1 (C)). This design enables the system to decompose the high-dimensional task into manageable sub-tasks, each assessing tabular data from different dimensions: (i) Planner Agent ($PA$) that designs the evaluation strategy. (ii) Specialized Agents ($SA$), $\{SA_{1},\dots,SA_{N}\}$, which are dynamically configured by the $PA$ to execute this strategy based on their role definition and expertise. (iii) Supervisor Agent ($SPA$) that aggregates analyses to make a final decision.
@@
Planner Agent ($PA$). The primary role of $PA$ is to interpret the high-level task context $\mathcal{Q}$ (Lead VC and target company profiles) and formulate an evaluation plan. It does so by analyzing $\mathcal{Q}$ with feature names to identify the most critical evaluation dimensions given the case description. The $PA$’s output is a set of $N$ agent configurations, $\{A_{1},\dots,A_{N}\}$, where each configuration $A_{i}$ contains a specific “profile” that guides the corresponding $SA_{i}$. It also generates “strategic guidance” to support the decision for the $SPA$. Formally, the output can be expressed as: $\{A_{1},\dots,A_{k}\},PA(\mathcal{Q},C_{sample})]$, a set of explicit instructions for the team of $SA$s.
@@
Specialized Agent ($SA$). Each $SA$ acts as a domain expert, tasked with evaluating the entire candidate pool $C$ from its specialized perspective. To manage the high dimensionality of the candidate feature vectors ($x_{j}\in\mathbb{R}^{n}$), the $SA$ first performs a feature selection step. Guided by its assigned $profile_{i}$, it identifies and focuses on a relevant feature subset. The agent’s core evaluation function $f_{i}$, driven by a backbone LLM, then directly produces a ranked shortlist of candidates. This output, $S^{\prime}_{i}$, is a list containing the top $k^{\prime}=\lceil m/3\rceil$ firms, where each entry includes the firm’s ID, its rank, and an alignment score, $score_{ij}\in[1,10]$.
@@
Supervisor Agent ($SPA$). The $SPA$ is responsible for synthesizing the shortlists $\{F^{\prime}_{1},\dots,F^{\prime}_{N}\}$ from Specialized Agents into the final ranked list $F$. It mimics a human-led committee’s decision-making process by first establishing consensus and then resolving disagreements based on the strategic priorities of the deal. This is achieved in a two-step process: (i) Consensus Selection: The $SPA$ first identifies candidates with broad support across $SA$, by counting how many agents include them in their shortlists. This step can determine robust candidates that perform well across different dimensions: $F_{1}(c_{j})=\sum_{i=1}^{N}\mathbb{I}(c_{j}\in F^{\prime}_{i})$. (ii) Conflict Resolution: To fill the remaining slots in the shortlist, the $SPA$ resolves disagreements among the $SA$s. It first determines an importance weight $w_{i}$ for each agent based on the agent’s relevance to the specific deal. It then resolves the conflict for the remaining candidates by giving more weight to the opinions of more important $SA$s. This allows the $SPA$ to select candidates who excel in critical areas, even if they lack broad consensus: $F_{2}(c_{j})=\sum_{i=1}^{N}w_{i}\cdot\frac{1}{R_{i}(c_{j})}$. The final shortlist $F=[F_{1}(c_{j}),F_{2}(c_{j})]$.